# Towards Effectively Detecting and Explaining Vulnerabilities Using Large Language Models

### v1

@@ 3.1. Vulnerability Interpretation Enhancement Prompting
Despite the strong code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning, especially in tasks that require a deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities. These challenges manifest as insufficient detection accuracy and vague vulnerability analysis. To guide LLMs towards better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques and the contextual learning capabilities of LLMs. By combining these with prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
@@
Figure 3 shows the design of our prompt templates. The prompt templates consist of four main components: task description, specific instructions, generation examples, and sample inputs. (1) Task description: provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format, helping the LLM understand task requirements and grasp the background knowledge of the vulnerabilities. (2) Specific instructions: include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output. These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis. (3) Generation examples: provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals. (4) Sample input: is used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
Figure 3. Prompting template of Automated Vulnerability Interpretation Labelling.
@@
Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, leading to well-performing fine-tuned specialized vulnerability detection and explanation models. This confirms the effectiveness of our prompt design approach.
3.2. Key Code Extraction Based on Chain-of-Thought (CoT)
@@
One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets. Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs, but these methods are difficult to apply in diverse vulnerability detection scenarios. To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction. This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code. The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities in the form of a thinking chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
@@
Figure 4 illustrates the prompt template for key statement extraction designed for LLMs. The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on. After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT. This guides the model in reference to the key statements step by step to complete the vulnerability detection, localization, and explanation.
Figure 4. Prompting template of Key Code Extraction.
4. Results
@@
4.1. RQ1: How effective are LLMs in detecting software vulnerabilities?
In this RQ, we discuss the detection performance of fine-tuned, vulnerability-specialized LLMs in various scenarios.
4.1.1. Experimental Setup.
Dataset. As discussed in Section 3, we select the SeVC (Li et al., 2021c) dataset and the DiverseVul dataset (Chen et al., 2023). assessments involving ten CWE types of vulnerabilities.
Backbone LLMs. We use the Codellama (Roziere et al., 2023) and Llama3 as our base models for fine-tuning. CodeLlama is initialized with the weights of Llama2 (Touvron et al., 2023) and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. Considering the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct (cod, 2023) as our primary model and used the 7B version and the newly released Llama3-8B-Instruct for comparison (refer to the discussion section).
Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability models, we design four sub-research questions:
@@ 4.1. RQ1: How effective are LLMs in detecting software vulnerabilities?
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenarios in RQ1.1 and RQ1.2, we directly use these three metrics. For the multi-class classification in RQ1.3, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.4, we use Micro-F1 and Macro-F1. To test the detection accuracy, we selected CodeT5 (Wang et al., 2021) and CodeBERT (Feng et al., 2020) as baselines and fine-tuned them for classification tasks by adding a linear classification layer.
Implementation Details. We implement our approach using the Transformers (Thapa et al., 2022) and PEFT (Mangrulkar et al., 2022) libraries with the PyTorch platform. All experiments are conducted on NVIDIA A100-SXM4-80GB GPU platforms, with the token length limit set to 2048. We use the AdamW optimizer and train the models for 3 epochs. For Low-rank Adaptation (LoRA) configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
4.1.2. RQ1.1: Detection with specified vulnerability type using a dedicated model.
To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type. Specifically, we indicate the type of vulnerability we are concerned with in the training and inference prompts to help it narrow down the scope of detecting vulnerabilities. During training, we use the explanations with vulnerability location information and specific details annotated by GPT as the target output for the model to perform detection and explanation tasks. The inputs for the model during training and inference are code snippets of the given examples. During inference, we develop a rule to analyze the generated result to determine whether it indicates a vulnerability or not. Specifically, we consider a result as non-vulnerable if it contains the sentences following a fixed pattern (e.g., “There is no security issues” and “detected in the given code snippet.”). Otherwise, we consider the result indicates a vulnerability. We randomly selected 100 vulnerable results and 100 non-vulnerable results to examine the correctness of such rule-based analysis. We find that all of the vulnerable and non-vulnerable results can be correctly classified by this rule.
Table 2. Performances of Single-Type Detection on SeVC(RQ1.1).
@@
Experimental Results. The experimental results are shown in Table 2. The fine-tuned models on four types of vulnerabilities have exceeded 90% performances on three metrics, which indicates that the specialized models for a single type of vulnerability can achieve good results. On one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type. On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs. If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
4.1.3. RQ1.2: Detection with specified vulnerability type using an all-in-one model.
To explore the impact of training with data from other types of vulnerabilities on the detection effect of the target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning on four types of vulnerabilities and performing individual detection for each type. Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection, and during inference, we also provide the vulnerability type information for each example. The difference between RQ1.1 and RQ1.2 is that, we fine-tune a dedicated model for each type of vulnerability in RQ1.1. In RQ1.2, we fine-tune an all-in-one model and use the prompts to distinguish the type of vulnerability in interest.
Table 3. Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
@@
Experimental Results. We present the results of the fine-tuned all-in-one model in Table 3. In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities. This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities. This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs. Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
4.1.4. RQ1.3: Detection with an identification of the vulnerability type.
In real software development environments, the vulnerability risks faced are often diverse. It may not always be possible to preemptively sense the potential types of vulnerabilities in the samples to be detected. This requires the large model to correctly identify the type of vulnerabilities present from multiple types and further provide specific explanations. To test the accuracy of a unified model in identifying and classifying code with multiple types of vulnerabilities, we conduct a multi-class vulnerability detection fine-tuning on SeVC. Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including none-vulnerable and the four types of vulnerabilities in SeVC. In the training and inference prompts, we no longer specify the type of vulnerability, but instead turn to a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
Table 4. Performances of Multi-Type Detection on SeVC(RQ1.3).
Precision Recall F1 Score
@@
Experimental Results. We present the metrics for each type and the confusion matrix for the multi-class vulnerability detection task in Table 4. Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code. (2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario. (3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types. The other three types exhibit a significant degree of confusion and are difficult to identify. This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient. The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
4.1.5. RQ1.4: Identification of the vulnerability type on the datasets with more vulnerability types.
@@
In real development environments, the security risks of vulnerabilities come from a greater variety of project types and more complex code structures, often involving a wider range of vulnerability models. To more closely resemble real development conditions, we conduct a multi-label classification task of the top 10 types of CWE on the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels). Specifically, we modify the task description to be based on CWE-type vulnerability detection and explanation tasks, and add the ‘[CWE]’ tag for generating a list of CWEs related to the target code. At the same time, we incorporate the CWE descriptions as part of the output explanation information to help the model further understand the meaning of each CWE type of vulnerability and to use the CWE descriptions for vulnerability analysis and identification.
Table 5. Performances of Multi-Label CWE Detection on DiverseVul(RQ1.4).
Metric CWE ID Overall
787 125 119 20 703 416 476 190 200 399 Micro Macro
Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 73.3%
Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 73.4%
F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 73.2%
@@
Experimental Results. The experimental results are presented in Table 5. From the results, we find that after fine-tuning, the specialized vulnerability models have achieved good detection accuracy for the 10 types of CWE, demonstrating an ability to grasp the pattern information for each CWE type. Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories. Each vulnerability type has a more fine-grained explicit definition, and it also avoids the potential category association issues that might exist in SeVC, thus resulting in better performance. This validates the feasibility of research into LLMs for vulnerability detection and explanation that are oriented toward real software development environments.
@@
4.2. RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities. Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanation. The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content. Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics, Accuracy, Clarity, and Constructiveness, as the evaluation criteria for vulnerability explanation. We use manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
4.2.1. Experimental Setup.
In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT. We randomly sample the data based on 95% confidence level and 5% confidence interval (Boslaugh and Watters, 2008) for the fine-tuned model’s correctly classified instances, obtaining 384 and 224 samples from SeVC and DiverseVul, respectively. Our evaluation includes two parts, (1) Manual review: two authors of this paper independently conduct manual checks on the samples, scoring them according to the criteria. In cases of disagreement, discussions are held until a consensus. The value of Cohen’s Kappa (McHugh, 2012) in this process is 0.76, which indicates a substantial agreement. (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion. We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction. This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.

### v2

@@ III-A Vulnerability Interpretation Enhancement Prompting
Despite the robust code understanding and analysis capabilities of LLMs, they still face challenges in complex reasoning tasks that require deep understanding of code, strong reasoning abilities, and specialized knowledge of vulnerabilities. These challenges manifest as insufficient detection accuracy and vague vulnerability analysis. To guide LLMs toward better understanding the goals of vulnerability explanation and to enhance their comprehension and analysis of vulnerable code, we have integrated instruction-based fine-tuning techniques with the contextual learning capabilities of LLMs. By leveraging prior knowledge of open-source code vulnerabilities, we design prompt templates for data annotation, fine-tuning, reasoning, and evaluation, effectively empowering the application of LLMs in vulnerability detection and explanation tasks across all crucial stages of the framework.
@@
Figure 3 illustrates the design of our prompt templates. These templates consist of four main components: task description, specific instructions, generation examples, and sample inputs. (1) Task description: Provides a specific template for the current vulnerability detection and explanation, including information on the types of vulnerabilities being detected and the basic input-output format. This helps the LLM understand task requirements and grasp the background knowledge of the vulnerabilities. (2) Specific instructions: Include requirements for the LLM’s input-output format, such as output steps, the range of vulnerability types to focus on, and the length of the output. These requirements leverage the LLM’s instruction-following ability to ensure uniform and standardized output formats, facilitating subsequent content use and analysis. (3) Generation examples: Provide manually screened samples of vulnerable code snippets and effective explanation data pairs to help the annotation model better understand the task and generation goals. (4) Sample input: Used for effectively inputting the code to be explained, and during the annotation stage, it also includes corresponding labels and other relevant information such as CWE, CVE descriptions, or commit messages as supplementary inputs.
Figure 3: Prompting template of Automated Vulnerability Interpretation Labelling.
@@
Through testing and evaluating various vulnerability detection and explanation tasks, we have verified that the prompt templates effectively achieved the expected goals in all stages, resulting in well-performing fine-tuned specialized vulnerability detection and explanation models.
III-B Key Code Extraction Based on Chain-of-Thought (CoT)
@@
One of the major challenges in explaining code vulnerabilities is extracting key code statements closely related to the target vulnerability type from long code snippets. Traditional static analysis methods extract code pattern information through abstract syntax trees, program dependence graphs, or control flow graphs. However, these methods are often difficult to apply in diverse vulnerability detection scenarios. To further enhance the ability of large models to analyze complex code structures, we propose a chain of thought (CoT) enhancement method based on key code extraction. This method leverages prompt engineering to use LLMs for automated key code extraction of vulnerable code. The extracted key code guides the fine-tuning model in locating suspicious vulnerabilities through a thought chain, thereby enabling targeted detection and explanation of the target vulnerabilities.
@@
The extracted key statements are primarily based on the semantic information of the code itself and the type of vulnerability being focused on. After obtaining the key statements for each vulnerability data instance, these key statements are integrated into the prompts used during the model fine-tuning phase in the form of CoT. This approach guides the model step by step to reference the key statements to complete the vulnerability detection, localization, and explanation.
IV Results
@@
IV-A RQ1: How effective are LLMs in detecting software vulnerabilities?
Accurate detection is the foundation for correct vulnerability explanation. In this research question (RQ), we first discuss the detection performance of the fine-tuned specialized vulnerability models across various scenarios.
IV-A1 Experimental Setup
Dataset. As discussed in Section III, we select the SeVC [21] dataset and the DiverseVul dataset [11]. Based on the characteristics of the two datasets, we constructed three different vulnerability detection tasks to evaluate the detection performance of LLMVulExp: binary classification vulnerability detection (SeVC), coarse-grained multi-class vulnerability detection (SeVC), and CWE type-based multi-label vulnerability detection (DiverseVul).
Backbone LLMs. We use the Codellama [39] and Llama3 as our backbone LLMs for fine-tuning. CodeLlama is initialized with the weights of Llama2 [40] and fine-tuned on a specialized code dataset, thus possessing strong code understanding and generation capabilities. The Llama-family LLMs have been widely used by prior studies related to LLMs for software engineering [41, 42, 43]. In light of the natural language instruction comprehension capability of the Instruct version and the training cost, we selected CodeLlama-13B-Instruct [44] as our primary model for the experiments discussed in this chapter. For comparison purposes, we also considered the 7B version and the newly released Llama3-8B-Instruct in the discussion section.
Experimental Setting. To comprehensively evaluate the fine-tuned vulnerability LLMs, we design four sub-research questions:
@@ IV-A1 Experimental Setup
To evaluate the accuracy of vulnerability detection, we use Precision, Recall, and F1-Score as our evaluation metrics. For the binary classification scenario in RQ1.1, we directly use these three metrics. For the multi-class classification in RQ1.2, we use Weighted-F1 and Macro-F1. For the multi-label classification in RQ1.3, we use Micro-F1 and Macro-F1. To test the detection accuracy, we selected CodeT5 [45] and CodeBERT [46] as baselines, which have been commonly used by prior studies in software engineering [47, 48]. In particular, we fine-tune them for classification tasks by adding a linear classification layer.
Implementation Details. We implement our approach using the Transformers [17] and PEFT [49] libraries with the PyTorch platform. All experiments are conducted on two NVIDIA A100-SXM4-80GB GPUs platforms, with the token length limit set to 2048. We use the AdamW optimizer and train the models for 3 epochs. For LoRA configuration, we set the learning rate to 0.0003, weight decay to 0.01, the LoRA rank is set to 16, the LoRA scaling factor to 16 and the dropout to 0.05.
IV-A2 RQ1.1: Detection with specified vulnerability type using a dedicated model
To evaluate the detection capability for a single type of vulnerability, we fine-tune each vulnerability type on SeVC to obtain a specialized model for that particular type. The sample size of each type is shown in Table II. In the training and inference prompts, we specify the type of vulnerability to help narrow the scope of detection. During training, we use explanations annotated by GPT-3.5, which include vulnerability location information and specific details, as the target output for the model to perform detection and explanation tasks. The inputs for the model during both training and inference are code snippets from the given examples. As a generative model, we conduct a binary classification task based on the semantics of the generated text. For vulnerability samples, the model generates corresponding explanatory information, while for non-vulnerability samples, it outputs a predefined fixed pattern (e.g., “There are no security issues”).
TABLE II: Performances Comparison of fine-tuning on SeVC(RQ1.1).
Metric API Arith. Pointer Array Average
#Samples 20,294 5,968 38,040 16,680 80,982
Precision (LLMVulExp) 91.6% 90.3% 93.7% 95.3% 92.7%
Precision (CodeLlama) 61.8% 64.5% 70.2% 66.2% 65.7%
Recall (LLMVulExp) 94.5% 93.6% 91.2% 89.7% 92.3%
Recall (CodeLlama) 26.3% 30.4% 50.2% 36.6% 35.9%
F1 (LLMVulExp) 93.0% 91.9% 92.4% 92.4% 92.4%
F1 (CodeLlama) 36.9% 41.4% 58.6% 47.2% 46.0%
TABLE III: Performances of Single-Type Detection on SeVC(RQ1.1).
@@
Experimental Results. We first compared the impact of fine-tuning on the detection accuracy of CodeLlama. As shown in Table II, CodeLlama-13B-Instruct lacks precise vulnerability identification capabilities. Fine-tuning significantly improves the detection accuracy. Therefore, in subsequent tasks, we no longer compare with the original CodeLlama. From Table III we can observe that our generative model still achieves comparable performance to other fine-tuned classification models. On the one hand, this demonstrates that our method can effectively enhance the model’s understanding of this type of vulnerability, capturing the key patterns for determining its type. On the other hand, it also reflects that the detection task for a single type of vulnerability is relatively less challenging for LLMs. If the focus of vulnerability types in practice is relatively concentrated, superior performance can be achieved by fine-tuning with data of the target type.
IV-A3 RQ1.2: Detection with specified vulnerability type using an all-in-one model
To explore the impact of training with data from different types of vulnerabilities on the detection effectiveness for target vulnerabilities, we conduct experiments on SeVC by uniformly fine-tuning a model on four types of vulnerabilities and then performing individual detection for each type. Specifically, during training, we use all data from the four types, employing the same prompt as in (RQ1.1) to indicate the vulnerability type of each example for binary classification detection. During inference, we also provide the vulnerability type information for each example. The difference between RQ1.1 and RQ1.2 is that in RQ1.1, we fine-tune a dedicated model for each type of vulnerability, whereas in RQ1.2, we fine-tune an all-in-one model and use prompts to distinguish the type of vulnerability of interest.
TABLE IV: Performances Comparison between Dedicated Mode and All-in-one Mode(RQ1.2).
@@
Experimental Results. We present the results of the fine-tuned all-in-one model in Table IV. In terms of the performance comparison between the dedicated mode and the all-in-one mode, when given the types of vulnerabilities to be detected in the code, the all-in-one mode consistently achieves significant improvements across various metrics for all types of vulnerabilities. This indicates that the model can learn more patterns related to code vulnerabilities from data that explains multiple types of vulnerabilities, thereby enhancing its general vulnerability detection capabilities. This also reflects that increasing the data volume and the richness of the training data can further improve the performance of vulnerability-specialized LLMs. Moreover, this improvement is due to the ability to avoid confusion and interference that may arise from different types of vulnerabilities when the target type of the code to be detected is specified.
IV-A4 RQ1.3: Detection with an identification of the vulnerability type
In real software development environments, the vulnerability risks faced are often diverse, making it difficult to preemptively sense the potential types of vulnerabilities in the samples to be detected. This requires the large model to accurately identify and classify multiple types of vulnerabilities and provide specific explanations. To test the accuracy of a unified model in this context, we conduct a multi-class vulnerability detection fine-tuning on SeVC with all samples. Specifically, we add a new ‘[type]’ tag to the model’s response output to classify the vulnerability code, including non-vulnerable and the four types of vulnerabilities in SeVC. In the training and inference prompts, we no longer specify the type of vulnerability. Instead, we adopt a multi-type vulnerability detection scenario and provide specific descriptions of the four types of vulnerabilities in the task description.
TABLE V: Performances of Multi-Type Detection on SeVC(RQ1.3).
Precision Recall F1
@@
Experimental Results. We present the metrics for each type for the multi-class vulnerability detection task in Table V. Based on the experimental results, we find that: (1) The non-vulnerability type code has a high precision and recall rate, indicating that the model still possesses strong vulnerability code identification capability under multi-type vulnerabilities and can effectively distinguish non-vulnerable code. (2) Compared to the detection results when the type of interest is provided, the overall performance has declined to varying degrees, indicating that the model finds it more challenging to differentiate between the types of vulnerabilities in a multi-type scenario. (3) Specifically, for each type of vulnerability, the Arithmetic Expression performs relatively better than other types. The other three types exhibit a significant degree of confusion and are difficult to identify. This analysis suggests that while the model can effectively identify non-vulnerable code, it struggles with differentiating between various types of vulnerabilities, especially when the data volume for certain types is insufficient. The results highlight the importance of balanced and representative training data for improving the model’s performance across all vulnerability types.
IV-A5 RQ1.4: Identification of the vulnerability type in datasets with more vulnerability types
@@
In real development environments, security risks stem from a wide variety of project types and complex code structures, often involving a broader range of vulnerability models. To better resemble real development conditions, we conduct a multi-label classification task on the top 10 types of CWE in the real-world vulnerability dataset DiverseVul (where a small number of codes have multiple CWE labels). Specifically, we modify the task description to focus on CWE-type vulnerability detection and explanation tasks, adding a ‘[CWE]’ tag to generate a list of CWEs related to the target code. Additionally, we incorporate CWE descriptions as part of the output explanation to help the model better understand the meaning of each CWE type and utilize these descriptions for vulnerability analysis and identification.
TABLE VI: Performances of Multi-Label CWE Detection on DiverseVul (RQ1.4).
Metric CWE ID Overall (Micro)
787 125 119 20 703 416 476 190 200 399 Ours CodeT5 CodeBERT
Precision 79.1% 70.9% 79.3% 68.1% 46.9% 82.1% 74.4% 80.4% 63.6% 88.2% 74.4% 70.8% 65.3%
Recall 83.3% 72.5% 75.8% 71.0% 53.6% 78.0% 56.9% 84.9% 75.0% 83.3% 75.2% 63.1% 61.9%
F1 Score 81.2% 71.7% 77.5% 69.5% 50.0% 80.0% 64.4% 82.6% 68.9% 85.7% 74.8% 67.0% 63.5%
@@
Experimental Results. The experimental results are presented in Table VI. From the results, we find that after fine-tuning, the specialized vulnerability models achieve good detection accuracy for the 10 types of CWE, demonstrating their ability to grasp the pattern information for each CWE type. Unlike the multi-class task in SeVC, the multi-label CWE task does not require the model to differentiate between categories. Each vulnerability type has a more fine-grained and explicit definition, avoiding potential category association issues in SeVC and resulting in better performance. This validates the feasibility of using LLMs for vulnerability detection and explanation in real software development environments, which is illustrated in the following explanation example.
1 Network::FilterStatus Context::onNetworkNewConnection() { 2 onCreate(root_context_id_); 3 if (!wasm_->onNewConnection_) { 4 return Network::FilterStatus::Continue; 5 } 6 if (wasm_->onNewConnection_(this, id_).u64_ == 0) { 7 return Network::FilterStatus::Continue; 8 } 9 return Network::FilterStatus::StopIteration; 10 } {mdframed} [style=mystyle] [label] This function is vulnerable.
[cwe] This function is related to [’CWE-476’]. CWE-476 NULL Pointer Dereference: A NULL pointer dereference occurs when the application dereferences a pointer that it expects to be valid, but is NULL, typically causing a crash or exit.
[location] if (!wasm_->onNewConnection_)
[explanation] The function is vulnerable to a NULL pointer dereference. It checks if ’wasm_->onNewConnection_’ is NULL but does not handle the case when it is NULL. If ’wasm_->onNewConnection_’ is NULL, the function continues its execution, which could lead to a NULL pointer dereference and cause a crash or exit. The vulnerability is related to CWE-476: NULL Pointer Dereference.
TABLE VII: Example of Generated Explanation on DiverseVul.
@@
IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
In this RQ, we explore the capability of the LLMVulExp in explaining the detected vulnerabilities. Although many evaluation metrics exist in the field of text generation by LLMs, there are no comprehensive criteria for the evaluation of vulnerability explanations. The explanatory information generated by the model is intended to help software developers identify and mitigate potential vulnerability risks, ensuring the accuracy of vulnerability analysis content. Additionally, it should be readable and concise, enabling developers to efficiently acquire the information and provide actionable suggestions for vulnerability remediation.
Considering the characteristics and requirements of the vulnerability explanation task, we propose three evaluation metrics: Accuracy, Clarity, and Actionability, as the evaluation criteria for vulnerability explanations. We use both manual review and automated evaluation by LLMs to assess the explanatory capabilities based on these criteria.
IV-B1 Experimental Setup
In our study, we examine the explanations generated by the fine-tuned model for vulnerability code on the SeVC and DiverseVul test sets, as well as the ground truth annotations based on GPT-3.5. Due to the large scale of the dataset we use and the high evaluation cost, we randomly sample the data based on a 95% confidence level and a 5% confidence interval [50] following prior studies [51, 52, 53] for the fine-tuned model’s correctly classified instances. We obtain 384 and 224 samples from SeVC and DiverseVul, respectively. This sampling method ensures the statistical validity and representativeness of our evaluation.
Our evaluation includes two parts: (1) Manual review: Two authors of this paper, both of whom have extensive research experience in software vulnerabilities and software development, independently conduct manual checks on the samples, scoring them according to the criteria. In cases of disagreement, discussions are held until a consensus is reached. The value of Cohen’s Kappa [54] in this process is 0.76, indicating substantial agreement. This rigorous manual review process ensures the reliability and accuracy of our evaluation. (2) Automated review using LLM: For the LLM automated review, we utilize GPT-3.5, which is instructed in the prompt with a detailed description of the vulnerability explanation evaluation task and each criterion. We ask the LLM to output the scores for three criteria, with 1 indicating satisfaction and 0 indicating non-satisfaction.
This dual approach of manual and automated evaluation allows us to assess the quality of the explanations provided by the fine-tuned model and explore the potential of using LLMs for automated evaluation tasks, which can be valuable for scaling the evaluation process in the future.